# Mass Translation and Reflectivity Device Solar Sail Momentum Management Using Predictive Control

### v2

@@ 4.4 Proposed MPC-Based Momentum Management
Fig. 8 presents a comparison of the simulation results of three MPC momentum management controllers using the same tuning. The first and second MPC controllers are based on the LQ-MPC policy in (21) (MPC Strategy 1) with and without the additional RCD threshold. The threshold serves as another tuning knob that reduces actuator usage until a sufficiently large input is required. When the MPC controller calculates an RCD input below the threshold, it is ignored; larger inputs are quantized using a PWM-based method as in Section 3.3 before being applied. The first controller, using PWM-quantized RCD inputs without any threshold, is shown by the blue line, while the second controller, which applies a $50\%$ threshold of the maximum RCD torque value, is represented by the red line. In other words, the second controller with RCD threshold forces $\tau^{\text{RCD}}_{b3}=0$ when $\tau^{\text{RCD}}_{b3}<0.5\cdot u^{\text{RCD}}_{\text{max}}$. The third controller, using the iterative MPC policy from equation (23) (MPC Strategy 2), is indicated by the purple line, with both the RCD threshold and PWM-quantization embedded into the fixed virtual input throughout the iterations.
@@
Fig. 8(a) shows the momentum management actuator inputs, where an ideal performance is to reduce the usage of momentum management actuators with minimal number of RCD on-off cycles, RCD “on” time, and AMT moving distance, while successfully managing RWs angular momentum. Fig. 8(b) shows the RW angular momentum in all axes, where the black dashed lines are the soft constraints, and the red dashed lines are the hard constraints based on the RW saturation limits. Figs. 8(c) and 8(d) show the solar sail attitude and angular velocity, which are the trajectories that the RWs aim to track to zero. Since the RW PID controller tracks the desired attitude satisfactorily and momentum management is performed successfully, the solar sail attitude and angular velocity plots look identical across the three MPC momentum management controllers. However, there is a significant discrepancy in RCD usage as well as a slight RWs angular momentum history difference between the three MPC momentum management controllers.
To better analyze the results, the simulation is dissected into 3 stages of operational condition, where the initial maneuver between time 030000-3000 sec (approximately an hour), the transient state between time 3000130003000-13000 sec (approximately $3$ hours), and the steady state between time 130001600013000-16000 sec (approximately an hour). The momentum management input and RWs angular momentum history of each stage are shown in Figs 910, and 11. Quantitative data for each MPC controller’s actuation usage in the respective operational stages are provided in Tables 23, and 4. AMT and RCDs actuation usage of Solar Cruiser’s state-of-the-art method are also presented as a comparison.
@@
In the initial stage (030000-3000 sec), as shown in Fig. 9 and Table 2, the RWs quickly accumulate angular momentum as they conduct the maneuver to track the desired attitude. At $t=100$ sec, the MPC controllers are activated, and all controllers successfully manage the RWs angular momentum while maintaining attitude tracking. The first MPC controller without an RCD threshold (blue line) utilizes $29$ RCD on-off cycles, while the second and third controllers (with RCD threshold and backwards-iterative MPC, respectively) reduce RCD usage significantly, requiring only 11 and 0 cycles, respectively. Compared to Solar Cruiser’s PID-based momentum management strategy, the actuator usage in RCD time and AMT travel distance are significantly lower during the initial trajectory tracking stage.
During the transient stage (3000130003000-13000 sec), as shown in Fig. 10 and Table 3, the AMT is moving towards its optimal position to counteract the constant disturbance, and the roll axis RW angular momentum stabilizes with RCD actuation. The backwards-iterative MPC controller (purple line) extends the transient stage and delays the onset of steady-state RCD actuation, which results in fewer RCD on-off cycles, but with a larger AMT travel distance. Specifically, the backwards-iterative MPC controller exhibits small AMT movements when entering the steady-state operation period. This can be due to the necessity of continuous RCD actuation in steady state to counteract the constant disturbance, which results in an imbalance of the input weights in the MPC cost function. Considering the trade-off between AMT usage and RCD actuation, fewer RCD cycles and RCD time are more preferable, as the RCD is a relatively new technology with uncertain lifespan due to issues such as overheating and being rated for a limited number of on-off cycles.
In the steady-state results of Fig. 11 and Table 4, all three MPC controllers converge to the same steady-state RCD cycle, and the AMT position stabilizes. The backwards-iterative MPC controller, however, continues to exhibit small AMT movements, which is not ideal. In this case, the second controller (non-iterative MPC with RCD threshold) provides the best performance, maintaining minimal AMT motion and RCD usage.
@@
MPC1 (non-iter, no Thr) 29 545 18.9097 18.7628
MPC1-Thr (non-iter, with $50\%$ Thr) 1 79 18.9108 18.7619
MPC2 (iter-back) 0 0 18.9674 18.7073
Solar Cruiser 1 700 67.5393 66.9156
Table 2: Initial maneuver momentum management actuator usage for time 030000-3000 sec.
@@
MPC1 (non-iter, no Thr) 100 3809 0.4576 0.4588
MPC1-Thr (non-iter, with $50\%$ Thr) 60 3257 0.4576 0.4587
MPC2 (iter-back) 2 116 0.5204 0.5196
Solar Cruiser 0 0 24.4849 24.3993
Table 3: Transient state momentum management actuator usage for time 3000130003000-13000 sec.
@@
MPC1 (non-iter, no Thr) 30 1533 0.0000 0.0000
MPC1-Thr (non-iter, with $50\%$ Thr) 30 1759 0.0000 0.0000
MPC2 (iter-back) 30 1962 0.0345 0.0349
Solar Cruiser 0.5 2400 0.8059 0.7942
Table 4: Steady-state momentum management actuator usage for time 130001600013000-16000 sec.
@@ 4.5 Discussion and Robustness Assessment
A comparison of $30000$ seconds (approximately $8$ hours) simulation between the Solar Cruiser PID-based momentum management method and the proposed MPC-based approach is provided in Table 5. Compared to the state-of-the-art momentum management strategy used on Solar Cruiser, the proposed MPC-based controller in Section 3 effectively manages momentum while significantly reducing actuator usage. The substantial reduction in total AMT travel distance demonstrates improved efficiency, which conserves power resources and enhances actuator longevity. Although the MPC-based approaches result in a larger number of RCD on-off cycles compared to Solar Cruiser’s method, the total RCD on time is similar.
@@
MPC1 (non-iter, no Thr) 299 14167 19.3674 19.2217
MPC1-Thr (non-iter, with $50\%$ Thr) 231 13916 9.3685 19.2208
MPC2 (iter-back) 172 11538 19.5286 19.2675
Solar Cruiser 2 13900 119.4941 118.7224
@@ 5 Conclusion
The novel model predictive control (MPC)-based momentum management strategies presented in this paper demonstrated the ability to handle the coupled nature of the solar sail’s dynamics in a practical manner while minimizing actuator usage. Simulation results showed the effectiveness of the proposed MPC-based momentum management strategies in the presence of a constant worst-case disturbance torque. The proposed MPC strategies feature significant improvement in reducing momentum management actuation usage in contrast to the state-of-the-art decoupled controller framework while preserving simplicity for onboard implementation. The iterative backwards-in-time MPC algorithm further captures the RCD’s on-off actuation and dead-band thresholds in the prediction model, and improved actuation efficiency when computation capacity is allowed.

### v3

@@ 5.4 Proposed MPC-Based Momentum Management
Figure 8 presents a comparison of the simulation results of three MPC momentum management controllers using the same tuning. The first and second MPC controllers are based on the LQ-MPC policy in (39) (MPC Strategy 1) with and without the additional RCD threshold. The threshold serves as another tuning knob that reduces actuator usage until a sufficiently large input is required. When the MPC controller calculates an RCD input below the threshold, it is ignored; larger inputs are quantized using a PWM-based method as in Section 4.3 before being applied. The first controller, using PWM-quantized RCD inputs without any threshold, is shown by the blue line, while the second controller, which applies a $50\%$ threshold of the maximum RCD torque value, is represented by the red line. In other words, the second controller with RCD threshold forces $\tau^{\text{RCD}}_{b3}=0$ when $\tau^{\text{RCD}}_{b3}<0.5\cdot u^{\text{RCD}}_{\text{max}}$. The third controller, using the iterative MPC policy from equation (44) (MPC Strategy 2), is indicated by the purple line, with both the RCD threshold and PWM-quantization embedded into the fixed virtual input throughout the iterations.
@@
Figure 8(a) shows the momentum management actuator inputs, where an ideal performance is to reduce the usage of momentum management actuators with minimal number of RCD on-off cycles, RCD “on” time, and AMT moving distance, while successfully managing RWs angular momentum. Figure 8(b) shows the RW angular momentum in all axes, where the black dashed lines are the soft constraints, and the red dashed lines are the hard constraints based on the RW saturation limits. There is a significant discrepancy in RCD usage as well as a slight RWs angular momentum history difference between the three MPC momentum management controllers.
To better analyze the results, the simulation is dissected into 3 stages of operational condition (approximately one hour each), where the initial maneuver between time 035000-3500 sec, the transient state between time 350075003500-7500 sec, and the steady state between time 7500110007500-11000 sec. The momentum management input and RWs angular momentum history of each stage are shown in Figs 910, and 11. Quantitative data for each MPC controller’s actuation usage in the respective operational stages are provided in Tables 23, and 4. AMT and RCDs actuation usage of Solar Cruiser’s state-of-the-art method are also presented as a comparison.
@@
In the initial stage (035000-3500 sec), as shown in Figure 9 and Table 2, the RWs quickly accumulate angular momentum as they conduct the maneuver to track the desired attitude. At $t=100$ sec, the MPC controllers are activated, and all controllers successfully manage the RWs angular momentum while maintaining attitude tracking. The first MPC controller without an RCD threshold (blue line) utilizes $34$ RCD on-off cycles, while the second and third controllers (with RCD threshold and backwards-iterative MPC, respectively) reduce RCD usage significantly, requiring 0 cycles. Compared to Solar Cruiser’s PID-based momentum management strategy, the actuator usage in RCD time and AMT travel distance are significantly lower during the initial trajectory tracking stage.
During the transient stage (350075003500-7500 sec), as shown in Figure 10 and Table 3, the AMT is moving towards its optimal position to counteract the constant disturbance, and the roll axis RW angular momentum stabilizes with RCD actuation. The backwards-iterative MPC controller (purple line) extends the transient stage and delays the onset of steady-state RCD actuation, which results in fewer RCD on-off cycles, but with a larger AMT travel distance. Specifically, the backwards-iterative MPC controller exhibits small AMT movements when entering the steady-state operation period. This can be due to the necessity of continuous RCD actuation in steady state to counteract the constant disturbance, which results in an imbalance of the input weights in the MPC cost function. Considering the trade-off between AMT usage and RCD actuation, fewer RCD cycles and RCD time are more preferable, as the RCD is a relatively new technology with uncertain lifespan due to issues such as overheating and being rated for a limited number of on-off cycles.
In the steady-state results of Figure 11 and Table 4, all three MPC controllers converge to the same steady-state RCD cycle, and the AMT position stabilizes. The backwards-iterative MPC controller, however, continues to exhibit small AMT movements, which is not ideal. In this case, the second controller (non-iterative MPC with RCD threshold) provides the best performance, maintaining minimal AMT motion and RCD usage.
Table 2: Initial maneuver momentum management actuator usage for time 035000-3500 sec.
@@
MPC1 34 515 14.5574 19.2945
MPC1-Thr 0 0 14.5528 19.2927
MPC2 0 0 14.5390 19.2883
Solar Cruiser 1 700 15.2928 67.0687
Table 3: Transient state momentum management actuator usage for time 350075003500-7500 sec.
@@
MPC1 40 1098 0.1043 0.1222
MPC1-Thr 12 653 0.1040 0.1444
MPC2 5 291 0.2445 4.0863
Solar Cruiser 0 0 0.3418 0.2006
Table 4: Steady-state momentum management actuator usage for time 7500110007500-11000 sec.
@@
MPC1 35 1286 0.0009 0.0078
MPC1-Thr 35 2031 0.0012 0.0147
MPC2 35 2094 0.0888 0.9647
Solar Cruiser 0.5 3400 0 0
@@ 5.5 Discussion and Robustness Assessment
A comparison of $30000$ seconds (approximately $8$ hours) simulation between the Solar Cruiser PID-based momentum management method and the proposed MPC-based approach is provided in Table 5. Compared to the state-of-the-art momentum management strategy used on Solar Cruiser, the proposed MPC-based controller in Section 4 effectively manages momentum while significantly reducing actuator usage. The substantial reduction in total AMT travel distance demonstrates improved efficiency, which conserves power resources and enhances actuator longevity. Although the MPC-based approaches result in a larger number of RCD on-off cycles compared to Solar Cruiser’s method, the total RCD on time is lower.
@@
MPC1 299 13884 14.6607 19.4408
MPC1-Thr 237 13884 14.6582 19.4551
MPC2 230 13589 14.9879 26.5160
Solar Cruiser 4 15100 66.6281 118.3571
@@ 6 Conclusion
The novel MPC-based momentum management strategies presented in this paper demonstrated the ability to handle the coupled nature of the solar sail’s dynamics in a practical manner while minimizing actuator usage and respecting the limited onboard computation resources. Simulation results showed the effectiveness of the proposed MPC-based momentum management strategies in the presence of a constant worst-case disturbance torque. The proposed MPC strategies feature significant improvement in reducing momentum management actuation usage in contrast to the state-of-the-art decoupled controller framework while preserving simplicity for onboard implementation. The iterative backwards-in-time MPC algorithm further captures the RCD’s on-off actuation and dead-band thresholds in the prediction model, and improved actuation efficiency when computation capacity is allowed.